OpenStudy (howard-wolowitz):

Medal for UR help:

OpenStudy (michele_laino):

here we have to write the function g(f(x)). What is the function g(f(x))?

OpenStudy (howard-wolowitz):

f is the f(x) on and g is the (x) one

OpenStudy (howard-wolowitz):

I think it wonts you to combine them

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

hint: \[\Large g\left( {f\left( x \right)} \right) = \sqrt {f\left( x \right) + 1} = ...?\]

OpenStudy (michele_laino):

you have to replace f(x) with its definition, or formula

OpenStudy (howard-wolowitz):

crap i forgot

OpenStudy (howard-wolowitz):

ok I got another fraction

OpenStudy (michele_laino):

hint: \[\Large g\left( {f\left( x \right)} \right) = \sqrt {f\left( x \right) + 1} = \sqrt {\frac{1}{{{x^2} - 1}} + 1} = ...?\]

OpenStudy (howard-wolowitz):

(x+2) and (x-1)

OpenStudy (michele_laino):

we have to simplify that expression, as below: \[\Large \begin{gathered} g\left( {f\left( x \right)} \right) = \sqrt {f\left( x \right) + 1} = \sqrt {\frac{1}{{{x^2} - 1}} + 1} = \hfill \\ = \sqrt {\frac{{{x^2}}}{{{x^2} - 1}}} \hfill \\ \end{gathered} \] Now the last radical exists, if the subsequent inequality is checked: \[\Large \frac{{{x^2}}}{{{x^2} - 1}} \geqslant 0\]

OpenStudy (michele_laino):

so you have to solve this inequality: \[\Large \frac{{{x^2}}}{{{x^2} - 1}} \geqslant 0\]

OpenStudy (michele_laino):

do you know how to solve that inequality?

OpenStudy (michele_laino):

since the numerator is always positive and it is zero at x=0, then we have to solve this equivalent inequality: \[\Large {x^2} - 1 > 0\]

OpenStudy (howard-wolowitz):

i can so it its : 

OpenStudy (michele_laino):

ok so we can write this: domain of g(f(x)) is the intersection between these two sets: real line and (-infinity, -1) union (1,+infinity) so we have:  so the requested domain is: \[\left( { - \infty , - 1} \right) \cup \left( {1, + \infty } \right)\]